OpenStudy (winsteria):

Which equation is a related equation for this equation? 23/a = 8 (23 over a equals 8 ) A. a = 23 x 8 B. a = 8/23 C. a = 23/8 D. a + 8 = 23 CAN SOMEONE PLEASE HELP ME I'LL GIVE YOU A MEDAL AND FAN

OpenStudy (anonymous):

c

OpenStudy (anonymous):

Answear is C, a=8/23. It is the same equation except its fliped around. (which still makes it same equation)

OpenStudy (anonymous):

What you need to do is first get a in the numerator. then divide by 8.\[23=8a\] \[\frac{ 23 }{ 8 }=\frac{ 8a }{ 8 }\]

OpenStudy (winsteria):

THANK YOU ALL SO MUCH. YOU ALL GET FAN.

OpenStudy (anonymous):

yayy